### v2

@@ Abstract
Game theory relies heavily on the availability of cardinal utility functions, but in fields such as matching markets, only ordinal preferences are typically elicited. The literature focuses on mechanisms with simple dominant strategies, but many real-world applications lack dominant strategies, making the intensity of preferences between outcomes important for determining strategies. Even though precise information about cardinal utilities is not available, some data about the likelihood of utility functions is often accessible. We propose to use Bayesian games to formalize uncertainty about the decision-makers’ utilities by viewing them as a collection of normal-form games where uncertainty about types persists throughout all game stages. Instead of searching for the Bayes-Nash equilibrium, we study how uncertainty in utilities is reflected in uncertainty of strategic play. To do this, we introduce a novel solution concept called $\alpha$-Rank-collections, which extends $\alpha$-Rank to Bayesian games. This allows us to analyze strategic play in, for example, non-strategyproof matching markets, for which appropriate solution concepts are currently lacking. $\alpha$-Rank-collections characterize the expected probability of encountering a certain strategy profile under replicator dynamics in the long run, rather than predicting a specific equilibrium strategy profile. We experimentally evaluate $\alpha$-Rank-collections using instances of the Boston mechanism, finding that our solution concept provides more nuanced predictions compared to Bayes-Nash equilibria. Additionally, we prove that $\alpha$-Rank-collections are invariant to positive affine transformations, a standard property for a solution concept, and are efficient to approximate.
@@ 1 Introduction
A vNM function specifies the preferences of an individual precisely, but it is usually difficult to elicit (Hauser and Urban, 1979). However, due to the precise information about a decision-maker’s preferences, knowledge of the vNM functions allow for precise solution concepts as well – the prevalent one being the Nash equilibrium (NE) (Nash et al., 1950). The NE is a fixed point in strategy space that no player unilaterally wants to deviate from, giving a single (possibly mixed) strategy profile as the solution.
@@
Strategy-proof mechanisms are rare, and most mechanisms used in the field do not possess this desirable property. For example, the Boston mechanism is widely used in school choice applications, but we lack a principled approach to studying strategic behavior in mechanisms that are not strategy-proof. Unfortunately, Nash equilibrium strategies for ordinal games (such as matching with preferences) are much less well understood. Dasgupta and Maskin (1986) were the first to analyze equilibrium in games with discontinuous preferences. Reny (2016) and Carmona and Podczeck (2016) discussed existence of Nash equilibria in ordinal games with ordered preferences. See Reny (2020) for an up-to-date survey. If an ordinal Nash equilibrium exists, it is a NE for every possible vNM utility function that follows the ordinal preferences of the players. This is a strong requirement that is often not satisfied, and in many ordinal games, ordinal Nash equilibrium does not exist. Importantly, we lose essential information about the intensity of preferences. However, even in applications such as school choice or kidney exchanges, some vNM utility functions are more likely than others.
@@
Our contribution is threefold: first, we use Bayesian games to model probabilistic information about the vNM functions of other players [^2] In ordinal mechanisms such as matching markets, these vNM functions obey ordinal preferences by the players. However, in contrast to auction theory, we do not need interpersonal utility comparisons. as a collection of normal-form games where players remain uncertain about their types until the ex-post stage. Second, we extend $\alpha$-Rank to Bayesian games. This combination provides a principled way to study strategic interaction in games, where this was not possible so far. For example, we can analyze strategies in matching markets that are not strategy-proof. Finally, we use the popular Boston mechanism to illustrate the new method but also apply it to a version of the Hawk-Dove game, where there is uncertainty about other players’ types. The study of the Boston mechanism draws on an experimental paper by Featherstone and Niederle (2016). It is particularly insightful because we do not have game-theoretical predictions unless we assume we know the vNM utility functions of all players exactly, which is unreasonable to believe. Thirdly, we provide two theoretical results that establish invariance to the choice of the vNM representation and stability towards small variations in the vNM functions itself. Overall, the paper extends the recent literature on the MCC solution concept to Bayesian games and provides a novel approach to analyze ordinal games.
@@ 2 Related Literature
The field of utility theory is concerned with answering what strategic entities are striving for (Fishburn, 1968). After a long history of thought, the vNM utility function has become standard for describing preferences in the presence of risky outcomes of different choices and in game theory. Von Neumann and Morgenstern showed that, if an individual’s preferences satisfy certain axioms of rational behavior, there exists a utility function that the individual aims to maximize (Von Neumann and Morgenstern, 1944). The vNM utility function is a cardinal utility function, i.e., utility is measured on an interval scale (Händler, 1984).
@@ A Bayesian game is specified by a tuple $(\mathcal{N},\mathcal{S},\mathcal{V},F,u)$ where
In a Bayesian game’s interim state, the players know their type (i.e., vNM function) but not those of the others. Although we use the same formalization, we interpret the prior distributions as uncertainty about participants’ specific vNM function. This resembles the ex-ante state of a game, where each player only has distributional information about its type. This distribution describes uncertainty in its type and that of others. Furthermore, we assume that a player’s strategy is not conditioned on its drawn type. This way, Bayesian games allow us to reason about the strategies considering the collection of normal-form games.
@@ 5.2 Two-Sided Matching Market
We build upon the game setting introduced for the experimental setup considered by Featherstone and Niederle (2016) in their laboratory study. The aligned preference environment is an instance of the school choice problem [^5] Also known as the student-courses or hospital-residents problem.. It consists of five students (3 Tops, 2 Averages) and three schools (Gold, Silver, Bronze). The Gold school has two seats, while the Silver and Bronze schools only have one seat, such that one student will remain unmatched. We denote the set of possible outcomes by $\mathcal{O}_{i}=\{G,S,B,U\}$, where $G$, $S$, $B$ stand for the Gold, Silver, and Bronze school respectively, and $U$ for being unmatched. All schools prefer the three Tops to the two Averages. Note that the students among these groups are ranked equally, and a lottery decides who gets a seat among these groups. For example, if the two Averages apply for the Silver school, a lottery decides who gets the seat. The schools are assumed to be not strategic, i.e., they always submit their true preferences. All students prefer the Gold to Silver to the Bronze school. All of these choices are preferred to being unmatched. Therefore, we require students to submit a full preference ordering over the schools, i.e., $\mathcal{S}_{i}=\{(G,S,B),(G,B,S),(S,G,B),(S,B,G),(B,G,S),(B,S,G)\}$ for each student $i$. This environment is asymmetric because Tops and Averages are treated differently and have different strategies.
@@ 6 Conclusion and Future Work
In many decision problems, specific vNM functions are not available; not even the players know their vNM function with certainty. This can be particularly challenging in matching markets where non-strategyproof mechanisms are used, as the intensity of preferences over outcomes can greatly impact strategic play. However, until now, there has been a lack of formal approaches to analyzing these types of games strategically. We use Bayesian games to model uncertainty in players’ vNM functions via a prior distribution and assume that uncertainty over players’ types persists until the ex-post stage. This achieves two things. First, the modeling approach offers the option to enrich ordinal, or generally unspecific, preferences by uncertain information about the strength or intensity of preferences that agents typically have over alternatives. Second, it shifts the focus of analysis towards determining the expected strategic behavior for a collection of normal-form games, where players are uncertain about both their own and others’ utilities.
Previous approaches that aim to solve Bayesian games, such as the Bayesian Nash equilibrium, rationalize strategies by taking the whole collection of normal-form games into account. Each player knows her type in the interim stage and rationalizes over other players’ types and strategies. In turn, they are collapsing strategic play into a fixed point in the ex-ante strategy space, which suffers from high sensitivity to modeling assumptions and intractability. Instead, we propose analyzing players’ behavior for each game instance separately, weighting it according to the prior distribution to obtain an expected strategic play that captures uncertainty in behavior across different games. We use $\alpha$-Rank to capture players’ long-term behavior in a single normal-form game, providing a unique characterization of strategies that emerge from finite population replicator dynamics that represents how likely agents are in a particular strategy profile. Consequently, we introduce $\alpha$-Rank-collections, a solution concept that combines the agents’ distributional information about the intensity of preferences with the expressive prowess of $\alpha$-Rank. Importantly, it satisfies several desirable properties. For example, we show that it is invariant to positive affine transformations so that it is independent of the choice of vNM functions’ representatives and is efficient to approximate. Furthermore, we demonstrate in two examples that it provides a more nuanced analysis of strategic behavior than equilibrium predictions. In contrast to ordinal Nash equilibria, existence is guaranteed.

### v3

@@ Abstract
Game theory relies heavily on the availability of cardinal utility functions, but in fields such as matching markets, only ordinal preferences are typically elicited. The literature focuses on mechanisms with simple dominant strategies, but many real-world applications lack dominant strategies, making the intensity of preferences between outcomes important for determining strategies. Even though precise information about cardinal utilities is not available, some data about the likelihood of utility functions is often accessible. We propose to use Bayesian games to formalize uncertainty about the decision-makers’ utilities by viewing them as a collection of normal-form games. Instead of searching for the Bayes-Nash equilibrium, we study how uncertainty in utilities is reflected in uncertainty of strategic play. To do this, we introduce a novel solution concept called $\alpha$-Rank-collections, which extends $\alpha$-Rank to Bayesian games. This allows us to analyze strategic play in, for example, non-strategyproof matching markets, for which appropriate solution concepts are currently lacking. $\alpha$-Rank-collections characterize the expected probability of encountering a certain strategy profile under replicator dynamics in the long run, rather than predicting a specific equilibrium strategy profile. We experimentally evaluate $\alpha$-Rank-collections using instances of the Boston mechanism, finding that our solution concept provides more nuanced predictions compared to Bayes-Nash equilibria. Additionally, we prove that $\alpha$-Rank-collections are invariant to positive affine transformations, a standard property for a solution concept, and are efficient to approximate.
@@ 1 Introduction
A vNM function specifies the preferences of an individual precisely, but it is usually difficult to elicit (Hauser and Urban, 1979). However, due to the precise information about a decision-maker’s preferences, knowledge of the vNM functions allow for precise solution concepts as well – the prevalent one being the Nash equilibrium (NE) (Nash et al., 1950). A NE is a fixed point in strategy space that no player unilaterally wants to deviate from, giving a single (possibly mixed) strategy profile as the solution.
@@
Strategy-proof mechanisms are rare, and most mechanisms used in the field do not possess this desirable property. For example, the Boston mechanism is widely used in school choice applications, but we lack a principled approach to studying strategic behavior in mechanisms that are not strategy-proof. Unfortunately, Nash equilibrium strategies for ordinal games (such as matching with preferences) are much less well understood. Dasgupta and Maskin (1986) were the first to analyze equilibrium in games with discontinuous preferences. Reny (2016) and Carmona and Podczeck (2016) discussed existence of Nash equilibria in ordinal games with ordered preferences. See Reny (2020) for an up-to-date survey. If an ordinal Nash equilibrium exists, it is a NE for every possible vNM utility function that follows the ordinal preferences of the players. This is a strong requirement that is often not satisfied, and in many ordinal games, ordinal Nash equilibrium does not exist. Importantly, we lose essential information about the intensity of preferences. However, in applications such as school choice or kidney exchanges, some vNM utility functions are more likely than others.
@@
Our contribution is threefold: first, we use Bayesian games to model probabilistic information about the vNM functions of other players [^2] In ordinal mechanisms such as matching markets, these vNM functions obey ordinal preferences by the players. However, in contrast to auction theory, we do not need interpersonal utility comparisons. as a collection of normal-form games. Second, we extend $\alpha$-Rank to Bayesian games. This combination provides a principled way to study strategic interaction in games, where this was not possible so far. For example, we can analyze strategies in matching markets that are not strategy-proof. Finally, we use the popular Boston mechanism to illustrate the new method but also apply it to a version of the Hawk-Dove game, where there is uncertainty about other players’ types. The study of the Boston mechanism draws on an experimental paper by Featherstone and Niederle (2016). It is particularly insightful because we do not have game-theoretical predictions unless we assume we know the vNM utility functions of all players exactly, which is unreasonable to believe. Thirdly, we provide two theoretical results that establish invariance to the choice of the vNM representation and stability towards small variations in the vNM functions itself. Overall, the paper extends the recent literature on the MCC solution concept to Bayesian games and provides a novel approach to analyze ordinal games.
@@ 2 Related Literature
The field of utility theory is concerned with answering what strategic entities are striving for (Fishburn, 1968). After a long history of thought, the vNM utility function has become standard for describing preferences in the presence of risky outcomes of different choices and in game theory. Von Neumann and Morgenstern showed that, if an individual’s preferences satisfy certain axioms of rational behavior, there exists a utility function such that the individual acts as if to maximize this utility (Von Neumann and Morgenstern, 1944). The vNM utility function is a cardinal utility function, i.e., utility is measured on an interval scale (Händler, 1984).
@@ A Bayesian game is specified by a tuple $(\mathcal{N},\mathcal{S},\mathcal{V},F,u)$ where
In a Bayesian game’s interim state, the players know their type (i.e., vNM function) but not those of the others. Although we use the same formalization, we interpret the prior distributions as uncertainty about participants’ specific vNM function. Given a draw from the prior, we obtain a normal-form game where strategic behavior can be predicted. The distribution over vNM functions dictates the likelihood of each normal-form game occurring, and predictions must be weighted according to the ex-ante uncertainty to provide an overall prediction. Thus, Bayesian games enable us to reason about strategies by considering a collection of possible normal-form games.
@@ 5.2 Two-Sided Matching Market
We build upon the game setting introduced for the experimental setup considered by Featherstone and Niederle (2016) in their laboratory study. The aligned preference environment is an instance of the school choice problem [^5] Also known as the student-courses or hospital-residents problem.. It consists of five students (3 Tops, 2 Averages) and three schools (Gold, Silver, Bronze). The Gold school has two seats, while the Silver and Bronze schools only have one seat, such that one student will remain unmatched. We denote the set of possible outcomes by $\mathcal{O}_{i}=\{G,S,B,U\}$, where $G$, $S$, $B$ stand for the Gold, Silver, and Bronze school respectively, and $U$ for being unmatched. All schools prefer the three Tops to the two Averages. Note that the students among these groups are ranked equally, and a lottery decides who gets a seat among these groups. For example, if the two Averages apply for the Silver school, a lottery decides who gets the seat. The schools are assumed to be not strategic, i.e., they always submit their true preferences. All students prefer the Gold to Silver to the Bronze school. All of these choices are preferred to being unmatched. Therefore, we require students to submit a full preference ordering over the schools, i.e., $\mathcal{S}_{i}=\{(G,S,B),(G,B,S),(S,G,B),(S,B,G),(B,G,S),(B,S,G)\}$ for each student $i$. This environment is asymmetric because Tops and Averages are treated differently by the schools.
@@ 6 Conclusion and Future Work
In many decision problems, specific vNM functions are not available; not even the players know their vNM function with certainty. This can be particularly challenging in matching markets where non-strategyproof mechanisms are used, as the intensity of preferences over outcomes can greatly impact strategic play. However, until now, there has been a lack of formal approaches to analyzing these types of games strategically. We use Bayesian games to model uncertainty in players’ vNM functions via a prior distribution. This achieves two things. First, the modeling approach offers the option to enrich ordinal, or generally unspecific, preferences by uncertain information about the strength or intensity of preferences that agents typically have over alternatives. Second, it shifts the focus of analysis towards determining the expected strategic behavior for a collection of normal-form games, where players know their own type and those of all others in the interim stage, and the game-theoretic prediction needs to consider strategic play given the ex-ante uncertainty.
Previous approaches that aim to solve Bayesian games, such as the Bayesian Nash equilibrium, rationalize strategies by considering the whole opponents’ prior distribution. Each player knows her type in the interim stage and rationalizes over other players’ types and strategies. In turn, they are collapsing strategic play into a fixed point in the ex-ante strategy space, which suffers from high sensitivity to modeling assumptions and intractability. Instead, we propose analyzing players’ behavior for each game instance separately, weighting it according to the prior distribution to obtain an expected strategic play that captures uncertainty in behavior across different games. We use $\alpha$-Rank to capture players’ long-term behavior in a single normal-form game, providing a unique characterization of strategies that emerge from finite population replicator dynamics that represents how likely agents are in a particular strategy profile. Consequently, we introduce $\alpha$-Rank-collections, a solution concept that combines the agents’ distributional information about the intensity of preferences with the expressive prowess of $\alpha$-Rank. Importantly, it satisfies several desirable properties. For example, we show that it is invariant to positive affine transformations so that it is independent of the choice of vNM functions’ representatives and is efficient to approximate. Furthermore, we demonstrate in two examples that it provides a more nuanced analysis of strategic behavior than equilibrium predictions. In contrast to ordinal Nash equilibria, existence is guaranteed.